OpenStudy (anonymous):

Solve the equation 12x^3+x^2-9/2 x-1=0 given that -1/2 is a root

OpenStudy (anonymous):

You must factor out (x-(-1/2)) from the equation. Does anyone know how to do this?

OpenStudy (anonymous):

factor out a \((2x+1)\)

OpenStudy (anonymous):

Or use Ruffini's rule. ( http://en.wikipedia.org/wiki/Ruffini's_rule)

OpenStudy (callisto):

-1/2 is a root => (2x+1) is a factor of 12x^3+x^2-9/2 x-1  12x^3+x^2-9/2 x-1 = (2x+1) (12x^2-5x-2) = (2x+1)(3x-2)(4x+1) Put it =0, you can solve it!